OpenStudy (anonymous):

If m<2=x+8 and m<3=2x+7, what is m<2? A. 30 B. 33 C. 57 D. 25

OpenStudy (anonymous):



OpenStudy (anonymous):

is there anything else that goes with the question?

OpenStudy (anonymous):

No, that is all they give me.

OpenStudy (goldeverything):

The awnser is A

OpenStudy (anonymous):

Could you show me how you got the answer?

OpenStudy (anonymous):

no its D

OpenStudy (goldeverything):

i started at m<3=2x+7

OpenStudy (anonymous):

since you already have one angle which is 90 and you know that all triangles equal to 180

OpenStudy (goldeverything):

true

OpenStudy (anonymous):

you would first subtract 180 from 90

OpenStudy (anonymous):

meaning that both angles equal to 90

OpenStudy (anonymous):

soo you would add the expressions of both angles and set them to equal to 90 then solve

OpenStudy (anonymous):

and the answer is 30

OpenStudy (anonymous):

Not D

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

The answer is A?

OpenStudy (anonymous):

i mean 25

OpenStudy (anonymous):

sorry

jimthompson5910 (jim_thompson5910):

how are you getting 30? m < 2 + m < 3 = 90 x+8 + 2x + 7 = 90 3x + 15 = 90 3x = 90 - 15 3x = 75 x = 75/3 x = 25 m < 2 = x+8 m < 2 = 25 + 8 m < 2 = 33

OpenStudy (anonymous):

well x is 25 then you would plug it back in

OpenStudy (anonymous):

so the answer is 33

OpenStudy (anonymous):

i forgot to do the second part sorry :)

OpenStudy (anonymous):

Thank you so much @jim_thompson5910. You're really saving my butt today!

jimthompson5910 (jim_thompson5910):

glad to be of help

OpenStudy (anonymous):

:)